3 kg of sliced peaches were mixed with 12 kg of sliced bananas which cost $3/kg to make fruit salad which costs $4/kg. What is the price per kg of sliced peaches?
Answer by stanbon(75887)

3 kg of sliced peaches were mixed with 12 kg of sliced bananas which cost $3/kg to make fruit salad which costs $4/kg. What is the price per kg of sliced peaches?
-------------------------
Equation:
value + value = value
3*p + 12*3 = (3+12)*4
3p + 36 = 60
3p = 24
p = $8/kg (price of sliced peaches)
-------------
Cheers,
Stan H.
------------